OpenStudy (anonymous):

two numbers have a difference of 20. Find the numbers if the sum of their squares is a minimum value. x-y=20 x^2 + y^2 =?

OpenStudy (zzr0ck3r):

is this for calc?

OpenStudy (anonymous):

functions

OpenStudy (zzr0ck3r):

im confused by the question, find the numbers if the sum of their squares is a min value? thats exactly what it says?

OpenStudy (anonymous):

yea it just means you have to complete the square to get the vertex, which is the two numbers

OpenStudy (zzr0ck3r):

sorry I messed up

OpenStudy (zzr0ck3r):

(y+20)^2 + y^2 = 2y^2 + 40y + 400 = 0 = y^2 + 20y + 200 = 0 (y+10)^2= 200 y = +- 10sqrt(100)-10

OpenStudy (zzr0ck3r):

errrrrr y = +- 10sqrt(2)-10 sorry

OpenStudy (zzr0ck3r):

I dont think you need to complete the square to get the virtex but I did it just in case you needed it

OpenStudy (anonymous):

thanks alot!

OpenStudy (zzr0ck3r):

I forget the virtex formula something like -a/2b?

OpenStudy (zzr0ck3r):

vertex*

OpenStudy (zzr0ck3r):

-b/2a so -20/2 = -10

OpenStudy (anonymous):

yeah the answer in the textbook was +/-10

OpenStudy (zzr0ck3r):

word yeah the +- from completeing the square i c